Which of the ordered pairs
(0, 0), (–2, 10), (–1, –5), (–3, 9)
are solutions for the equation y = 5x?
No book. Thank you very much for your time!!

Answer by tutorcecilia(2152)

y = 5x [Basic equation]
Plug each ordered pair into the basic equation and solve to see if it is a true statement.
.
y = 5x [Plug in the ordered pair (0,0)
0=5(0)
0=0 [Correct -- this is a true statement]
.
y = 5x [Plug-in the ordered pair (-2, 10)
-10=5(-2)
-10=-10 [Correct--this is a true statement]
.
.
y = 5x [Plug-in the ordered pair (–1, –5)
-5=5(-1)
-5=-5 [Correct--a true statement]
.
.
y = 5x [Plug-in the ordered pair (–3, 9)]
9=5(-3)
9=-15 [Incorrect--this is an untrue statement, so (-3, 9) is not a solution to y=5=5x]